Prove that 1·1 ! + 2·2!+· ··+ n · n! = (n+I)!-1


whenever n is a positive integer.


1
Expert's answer
2022-02-02T16:17:55-0500

"\begin{aligned}
&\text { To proof: } 1 \cdot 1 !+2 \cdot 2 !+\ldots+n \cdot n !=(n+1) !-1 \text { for every positive integer } n . \\
&\text { PROOF BY INDUCTION: } \\
&\text { Let } P(n) \text { be } 1 \cdot 1 !+2 \cdot 2 !+\ldots+n \cdot n !=(n+1) !-1 \\
&\text { Basis step } n=1 \\
&1 \cdot 1 !+2 \cdot 2 !+\ldots+n \cdot n !=1 \cdot 1 !=1 \cdot 1=1 \\
&(n+1) !-1=(1+1) !-1=2 !-1=2-1=1 \\
&\text { We then note } P(1) \text { is true. } \\
&\text { Induction step: Let } P(k) \text { be true. } \\
&1 \cdot 1 !+2 \cdot 2 !+\ldots+k \cdot k !=(k+1) !-1 \\
&\text { We need to prove that } P(k+1) \text { is also true. }
\end{aligned}"

"\begin{aligned}
&1 \cdot 1 !+2 \cdot 2 !+\ldots+k \cdot k !+(k+1) \cdot(k+1) ! \\
&=(k+1) !-1+(k+1) \cdot(k+1) ! \\
&=1 \cdot(k+1) !+(k+1) \cdot(k+1) !-1 \\
&=(1+k+1)(k+1) !-1 \\
&=(k+2)(k+1) !-1 \\
&=(k+2) !-1 \\
&=((k+1)+1) !-1
\end{aligned}"

We then note that "P(k+1)" is also true.

Conclusion By the principle of mathematical induction, "P(n)" is true for all positive integers "n" .

Thus, "1\u00b71 ! + 2\u00b72!+\u00b7 \u00b7\u00b7+ n \u00b7 n! = (n+1)!-1" whenever n is a positive integer.

Hence Proved